OpenStudy (anonymous):

Simplify: sqrt[3]{81x ^{7}y ^{12}z ^{14}}

OpenStudy (anonymous):

\[\sqrt[3]{81x ^{7}y ^{12}z ^{14}}\]

OpenStudy (anonymous):

(81)^1/3=(3^3)1/3=3 ;x^7]^1/3=x^7/3 [y^12]^1/3=y^4[ [z^14]^1/3=z^14/3 3*(x^7/3)*(y^4)*(z^14/3)

OpenStudy (anonymous):

You have to use the fact that\[ \sqrt[b]{x^a} = x^{a/b} \]